OpenStudy (anonymous):

i need help please :D A ball has a diameter of 9 in. It consists of 2 parts. The inside is a spherical core with a diameter of 6 in. Surrounding the core is a layer of polyurethane. What is the volume of the polyurethane? Use 3.14 to approximate pi and express your answer in hundredths. ____in3

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@horsegirl325

OpenStudy (freckles):

find the volume of the whole ball then subtract out the volume of the spherical core

OpenStudy (anonymous):

oh okay. thanks

OpenStudy (freckles):

do you know the volume of a sphere?

OpenStudy (anonymous):

no D:

OpenStudy (freckles):

i was thinking you were going to say yes. it is weird your homework ask you to apply the volume of a sphere without you not having it in your notes.

OpenStudy (freckles):

\[V=\frac{4}{3} \pi r^3\] is the volume of a sphere this is the formula you will use to find the volume of the whole ball and also the formula you will use to find the volume of the spherical core r means radius you are given diameter you divide diameter in half to find the length of radius

OpenStudy (anonymous):

ohhh, okay, ill tell you what i get c:

OpenStudy (anonymous):

684*3.14 = 2147.76 in³. i got 2147.76

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

hmm how did you get that

OpenStudy (freckles):

did you find the difference yet of the two volumes ?

OpenStudy (anonymous):

the volume of the whole sphere is 972 the volume of the inner core is 288 972π - 288π = 684π

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

can i ask how you got the volume for the whole sphere ?

OpenStudy (freckles):

\[V_\text{ whole sphere}=\frac{4}{3} \cdot 3.14 \cdot (\frac{9}{2})^3\] all I did is use the formula above

OpenStudy (freckles):

that one should equal 381.51 try the second sphere ( the inner sphere) again